mac and his brother ted collect coins. Macs collection contains 8 more than twice as many coins as teds. if the total number of coins in the two collection is 74 how many coins are in macs collection
Answer by chessace(471)

m=8+2t
m+t=74
m=8+2(74-m)
3m=8+148=156
m=52